# Elusive groups from non-split extensions

### v1

@@ 4. Elusive groups of degree not divisible by 44
$(M\cap U)^{(1,4,5)}$ $=\langle u_{1}u_{2}u_{4},u_{2}u_{3}^{-1},u_{1}u_{3}\rangle=\langle u_{1}u_{2},u_{4}\rangle\times\langle u_{1}u_{2}^{-1}u_{3}^{-1}\rangle,$ (13)
$(M\cap U)^{(2,4,5)}$ $=\langle u_{1}u_{2}u_{4},u_{3}^{-1}u_{4},u_{2}u_{3}\rangle=\langle u_{1},u_{2}u_{4}\rangle\times\langle u_{1}u_{2}^{-1}u_{3}^{-1}\rangle,$ (14)
$(M\cap U)^{(3,4,5)}$ $=\langle u_{1}u_{2}u_{3},u_{2}u_{4}^{-1},u_{3}u_{4}\rangle=\langle u_{1},u_{2}u_{4}^{-1}\rangle\times\langle u_{1}u_{2}^{-1}u_{3}^{-1}\rangle,$ (15)
$(M\cap U)^{(4,2,5)}$ $=\langle u_{1}u_{4},u_{3}^{-1},u_{2}u_{3}u_{4}\rangle=\langle u_{1}u_{4},u_{2}u_{4}\rangle\times\langle u_{1}u_{2}^{-1}u_{3}^{-1}\rangle,$ (16)
$(M\cap U)^{(4,3,5)}$ $=\langle u_{1}u_{2}u_{4},u_{2},u_{3}u_{4}\rangle=\langle u_{1}u_{4},u_{2}\rangle\times\langle u_{1}u_{2}^{-1}u_{3}^{-1}\rangle.$ (17)
Denote the groups in (12)–(17) by $U_{1},\dots,U_{6}$ in order. The intersection of any two among $U_{1},\dots,U_{6}$ has order $9$, and the intersection of any four among $U_{1},\dots,U_{6}$ has order $3$. Moreover, for a $3$-subset $\{i,j,k\}$ of $\{1,\ldots,6\}$, we can directly verify that
$|U_{i}\cap U_{j}\cap U_{k}|=\begin{cases}9&\text{if }\{i,j,k\}=\{1,3,5\},\,\{1,4,6\},\,\{2,3,6\}\text{ or }\{2,4,5\}\\ 3&\text{else}.\end{cases}$
@@
$|U_{1}\cup\dots\cup U_{6}|=27\binom{6}{1}-9\binom{6}{2}+9\cdot 4+3\left(\binom{6}{3}-4\right)-3\binom{6}{4}+3\binom{6}{5}-3\binom{6}{6}=81.$
Therefore, the images of $M\cap U$ under elements in $G$ cover $U$, or equivalently, each element of $U$ is mapped into $M\cap U$ by some element of $G$. This proves that the transitive permutation group $X=(\mathrm{C}_{3}^{4}\times\mathrm{C}_{5}^{3}){}^{\boldsymbol{\cdot}}\mathrm{A}_{5}$ with stabilizer $Y=(\mathrm{C}_{3}^{3}\times\mathrm{C}_{5}^{2}){:}\mathrm{C}_{2}^{2}$ is elusive.
Now take a subgroup $W=M{:}\mathrm{C}_{2}$ of $Y$. Since there is only one conjugacy class of involutions in $G=\mathrm{A}_{5}$, the transitive permutation group $G$ with stabilizer $\overline{W}$ is $2$-elusive. Thereby we derive from Lemma 2.1 that the transitive permutation group $X$ with stabilizer $W$ is $2$-elusive. Moreover, our proof shows that $X$ is both $3$-elusive and $5$-elusive. Hence the transitive permutation group $X=(\mathrm{C}_{3}^{4}\times\mathrm{C}_{5}^{3}){}^{\boldsymbol{\cdot}}\mathrm{A}_{5}$ with stabilizer $W=(\mathrm{C}_{3}^{3}\times\mathrm{C}_{5}^{2}){:}\mathrm{C}_{2}$ is elusive, which completes the proof. ∎
@@ 5. Concluding remarks
An equivalent condition to (a) is that the induced permutation group of $E$ on each orbit is regular. If condition (a) holds, then $E_{u}=E_{v}$ for $u$ and $v$ in the same orbit of $E$. Also, under condition (a), $|E_{u}|=|E_{v}|$ for each $u$ and $v$ in $\Omega$ if and only if $G$ is half-transitive (all orbits have the same length). In the conclusion of the theorem, the existence of a derangement of prime order is equivalent to the existence of a derangement that is a product of disjoint cycles of the same length.
@@
Classify elusive groups $X=N.G$ with stabilizer $Y$ such that $G$ is a quasiprimitive $p^{\prime}$-elusive permutation group with stabilizer $YN/N$ for some prime $p$ and $N$ is an irreducible $\mathbb{F}_{p}G$-module.
@@ Acknowledgments
Chen was supported by the National Natural Science Foundation of China (12101518) and by the Fundamental Research Funds for the Central Universities (20720210036, 20720240136). O’Brien was supported by the Marsden Fund of New Zealand Grant 23-UOA-080 and by a Research Award of the Alexander von Humboldt Foundation.

### v2

@@ 4. Elusive groups of degree not divisible by 44
$(M\cap J)^{(1,4,5)}$ $=\langle w_{1}w_{2}u_{4},w_{2}w_{3}^{-1},w_{1}u_{3}\rangle=\langle w_{1}w_{2},w_{4}\rangle\times\langle w_{1}w_{2}^{-1}w_{3}^{-1}\rangle,$ (14)
$(M\cap J)^{(2,4,5)}$ $=\langle w_{1}w_{2}u_{4},w_{3}^{-1}u_{4},w_{2}w_{3}\rangle=\langle w_{1},w_{2}w_{4}\rangle\times\langle w_{1}w_{2}^{-1}w_{3}^{-1}\rangle,$ (15)
$(M\cap J)^{(3,4,5)}$ $=\langle w_{1}w_{2}w_{3},w_{2}w_{4}^{-1},w_{3}u_{4}\rangle=\langle w_{1},w_{2}w_{4}^{-1}\rangle\times\langle w_{1}w_{2}^{-1}w_{3}^{-1}\rangle,$ (16)
$(M\cap J)^{(4,2,5)}$ $=\langle w_{1}w_{4},w_{3}^{-1},w_{2}w_{3}w_{4}\rangle=\langle w_{1}w_{4},w_{2}w_{4}\rangle\times\langle w_{1}w_{2}^{-1}w_{3}^{-1}\rangle,$ (17)
$(M\cap J)^{(4,3,5)}$ $=\langle w_{1}w_{2}w_{4},w_{2},w_{3}w_{4}\rangle=\langle w_{1}w_{4},w_{2}\rangle\times\langle w_{1}w_{2}^{-1}w_{3}^{-1}\rangle.$ (18)
Denote the groups in (13)–(18) by $W_{1},\dots,W_{6}$ in order. The intersection of any two among $W_{1},\dots,W_{6}$ has order $9$, and the intersection of any four among $W_{1},\dots,W_{6}$ has order $3$. Moreover, for a $3$-subset $\{i,j,k\}$ of $\{1,\ldots,6\}$, we can directly verify that
$|W_{i}\cap W_{j}\cap W_{k}|=\begin{cases}9&\text{if }\{i,j,k\}=\{1,3,5\},\,\{1,4,6\},\,\{2,3,6\}\text{ or }\{2,4,5\}\\ 3&\text{else}.\end{cases}$
@@
$|W_{1}\cup\dots\cup W_{6}|=27\binom{6}{1}-9\binom{6}{2}+9\cdot 4+3\left(\binom{6}{3}-4\right)-3\binom{6}{4}+3\binom{6}{5}-3\binom{6}{6}=81.$
Therefore, the images of $M\cap J$ under elements in $G$ cover $J$, or equivalently, each element of $J$ is mapped into $M\cap J$ by some element of $G$. This proves that the transitive permutation group $X=(\mathrm{C}_{3^{k}}^{4}\times\mathrm{C}_{5}^{3}){}^{\boldsymbol{\cdot}}\mathrm{A}_{5}$ with stabilizer $Y=(\mathrm{C}_{3^{k}}^{3}\times\mathrm{C}_{5}^{2}){:}\mathrm{C}_{2}^{2}$ is elusive.
Note that $\overline{W}$ contains an involution $\overline{x}$. Since there is only one conjugacy class of involutions in $G=\mathrm{A}_{5}$, the transitive permutation group $G$ with stabilizer $\overline{W}$ is $2$-elusive. Thereby we derive from Lemma 2.1 that the transitive permutation group $X$ with stabilizer $W$ is $2$-elusive. Moreover, our proof shows that $X$ is both $3$-elusive and $5$-elusive. Hence the transitive permutation group $X=(\mathrm{C}_{3^{k}}^{4}\times\mathrm{C}_{5}^{3}){}^{\boldsymbol{\cdot}}\mathrm{A}_{5}$ with stabilizer $W=(\mathrm{C}_{3^{k}}^{3}\times\mathrm{C}_{5}^{2}){:}\mathrm{C}_{2}$ is elusive, which completes the proof. ∎
Remark.
As previously explained, the presentation of the group $X$ was derived from the action of $\mathrm{A}_{5}$ on its fully deleted permutation modules in characteristic $3$ and $5$. In the proof of Theorem 1.4, we were able to use relations coming from more general modules for $\mathrm{A}_{5}$ over $\mathrm{C}_{3^{k}}$ with $k\geq 1$ an integer. However, the following shows that the same approach does not apply in the case of characteristic $5$.
Fix an integer $\ell\geq 1$. Let $X_{\ell}$ be a group with generators $u_{1},u_{2},u_{3},u_{4},v_{1},v_{2},v_{3},x,y$ satisfying the relations in the presentation of the group $X$ from Construction 4.2 with the relation $v_{i}^{5}=1$ replaced by $v_{i}^{5^{\ell}}=1$. Note that $V\coloneqq\langle v_{1},v_{2},v_{3}\rangle$ is an abelian normal subgroup of $X_{\ell}$ and a quotient of $\mathrm{C}_{5^{\ell}}^{3}$. Since $(xy)^{5}$ and $v_{1}$ are in $V$,
$v_{1}^{(xy)^{5}}=v_{1}.$
The left-hand side can be calculated directly as follows:
$v_{1}^{xy}$ $=v_{2}^{y}=v_{1}^{-1}v_{2},$
$v_{1}^{(xy)^{2}}$ $=(v_{1}^{-1}v_{2})^{xy}=(v_{2}^{-1}v_{1})^{y}=(v_{1}^{-1}v_{2})^{-1}(v_{1}^{-1}v_{3})=v_{2}^{-1}v_{3},$
$v_{1}^{(xy)^{3}}$ $=(v_{2}^{-1}v_{3})^{xy}=(v_{1}^{-2}v_{2}^{-1}v_{3}^{-1})^{y}=(v_{1}^{-1}v_{3})^{-2}(v_{1}^{-1}v_{2})^{-1}v_{1}=v_{1}^{4}v_{2}^{-1}v_{3}^{-2},$
$v_{1}^{(xy)^{4}}$ $=(v_{1}^{4}v_{2}^{-1}v_{3}^{-2})^{xy}=(v_{1}v_{2}^{6}v_{3}^{2})^{y}=(v_{1}^{-1}v_{3})(v_{1}^{-1}v_{2})^{6}(v_{1}^{-1})^{2}=v_{1}^{-9}v_{2}^{6}v_{3},$
$v_{1}^{(xy)^{5}}$ $=(v_{1}^{-9}v_{2}^{6}v_{3})^{xy}=(v_{1}^{5}v_{2}^{-10}v_{3}^{-1})^{y}=v_{1}^{6}v_{2}^{-10}v_{3}^{5}.$
Hence we derive from $v_{1}^{(xy)^{5}}=v_{1}$ that
$v_{1}^{5}v_{2}^{-10}v_{3}^{5}=1.$ (19)
Note that $(yx)^{5}=y(xy)^{5}y^{-1}=((xy)^{5})^{y^{-1}}$. Since $(xy)^{5}\in V$ and $V$ is normal in $X_{\ell}$, we conclude that $(yx)^{5}\in V$, and so
$v_{2}^{(yx)^{5}}=v_{2}.$
The left-hand side can be calculated as follows by noting $(xy)^{5}=x(yx)^{5}x$:
$v_{2}^{(yx)^{5}}=v_{2}^{x(yx)^{5}x}=\left(v_{1}^{(xy)^{5}}\right)^{x}=(v_{1}^{6}v_{2}^{-10}v_{3}^{5})^{x}=v_{1}^{-15}v_{2}v_{3}^{-5}.$
Thus, we derive from $v_{2}^{(yx)^{5}}=v_{2}$ that
$v_{1}^{15}v_{3}^{5}=1.$ (20)
Moreover, $v_{3}$ and $(yx)^{5}$ also commute, implying that
$v_{3}^{(yx)^{5}}=v_{3}.$
Since $y^{3}\in U\coloneqq\langle u_{1},u_{2},u_{3},u_{4}\rangle$, we deduce that $v^{y^{3}}=v$ and hence $v^{y^{-1}}=v^{y^{2}}$ for all $v\in V$. Therefore,
$v_{3}^{(yx)^{5}}=v_{3}^{y(xy)^{5}y^{-1}}=\left(\left(v_{1}^{(xy)^{5}}\right)^{-1}\right)^{y^{2}}=(v_{1}^{-6}v_{2}^{10}v_{3}^{-5})^{y^{2}}=(v_{1}v_{2}^{10}v_{3}^{-6})^{y}=v_{1}^{-5}v_{2}^{10}v_{3}.$
We conclude that
$v_{1}^{-5}v_{2}^{10}=1.$ (21)
By multiplying (19) and (21), we obtain $v_{3}^{5}=1$. This combined with (20) implies that $v_{1}^{15}=1$, which is equivalent to $v_{1}^{5}=1$ (since $V$ is a quotient of a $5$-group). Plugging this back into (21) gives $v_{2}^{10}=1$, further implying that $v_{2}^{5}=1$. Hence, $X_{\ell}$ is isomorphic to $X$ for every integer $\ell\geq 1$.
@@ 5. Concluding remarks
An equivalent condition to (a) is that the induced permutation group of $E$ on each orbit is regular. If condition (a) holds, then $E_{u}=E_{v}$ for $u$ and $v$ in the same orbit of $E$. Also, under condition (a), $|E_{u}|=|E_{v}|$ for each $u$ and $v$ in $\Omega$ if and only if $E$ is half-transitive (all orbits have the same length). In the conclusion of the theorem, the existence of a derangement of prime order is equivalent to the existence of a derangement that is a product of disjoint cycles of the same length.
@@
Classify elusive groups $N.G$ with stabilizer $Y$ such that $G$ is a quasiprimitive $p^{\prime}$-elusive permutation group with stabilizer $YN/N$ for some prime $p$ and $N$ is an irreducible $\mathbb{F}_{p}G$-module. In particular, is there any counterexample to the Polycirculant Conjecture among such groups $X$?
@@ Acknowledgments
We thank the referee for careful and insightful review of our manuscript. Chen was supported by the National Natural Science Foundation of China (12101518) and by the Fundamental Research Funds for the Central Universities (20720210036, 20720240136). Lee was supported by an Australian Research Council Discovery Early Career Researcher Award (project number DE230100579). Part of the work was done during Mitrović’s visit to Lee at Monash University and Xia at the University of Melbourne, supported by the Marsden Fund of New Zealand Grant 21-UOA-174 and University of Auckland’s DRDF Grant. O’Brien was supported by the Marsden Fund of New Zealand Grant 23-UOA-080 and by a Research Award of the Alexander von Humboldt Foundation. Xia was supported by ARC Discovery Project DP250104965.